Apisco Tiger Inc. has bonds outstanding with original maturity of 13 years (at the time of issue). The bonds have a 7.42% annual coupon rate (coupons paid annually) and were issued 4 years ago at their par value of $1,000. However, due to changes in interest rates, the bond's market price has increased to $1,090.83. What is the bond's yield to maturity today?

A. 6.08 percent

B. 6.37 percent

C. 7.56 percent

D. 6.57 percent
Answer #1

YTM: It is the discount rate at which present value of a bond’s
promised cash flows is equal to its market price is the bond.

YTM : 6.0799%

Answer A. 6.08 percent.

original maturity - 13 years. were issued 4 years ago. So remaining maturity= 13-4e a years. AT YTM 74-2 1090.83 = 1000 (144m